# Comparative Study of Causal Discovery Methods for Cyclic Models with Hidden Confounders

### v1

@@ Abstract
Nowadays, the need for causal discovery is ubiquitous. A better understanding of not just the stochastic dependencies between parts of a system, but also the actual cause-effect relations, is essential for all parts of science. Thus, the need for reliable methods to detect causal directions is growing constantly. In the last 50 years, many causal discovery algorithms have emerged, but most of them are applicable only under the assumption that the systems have no feedback loops and that they are causally sufficient, i.e. that there are no unmeasured subsystems that can affect multiple measured variables. This is unfortunate since those restrictions can often not be presumed in practice. Feedback is an integral feature of many processes, and real-world systems are rarely completely isolated and fully measured. Fortunately, in recent years, several techniques, that can cope with cyclic, causally insufficient systems, have been developed. And with multiple methods available, a practical application of those algorithms now requires knowledge of the respective strengths and weaknesses. Here, we focus on the problem of causal discovery for sparse linear models which are allowed to have cycles and hidden confounders. We have prepared a comprehensive and thorough comparative study of four causal discovery techniques: two versions of the LLC method [10] and two variants of the ASP-based algorithm [11]. The evaluation investigates the performance of those techniques for various experiments with multiple interventional setups and different dataset sizes.
@@ I Introduction
Causal analysis [15, 16] of complex systems is nowadays an integral part of many sciences. It is heavily used in fields as diverse as medicine, biology, cognitive science, economics, predictive maintenance, root cause analysis, physics, and machine learning. Conventional data analysis investigates the probabilistic properties of the data to gain insight into the involved probability distributions which can then be used to e.g. predict new data. Causality on the other hand serves to not just learn the data but to learn about the system that generates the data. For instance, consider two random variables, one is binary and indicates the administration of a drug that is allegedly reducing blood pressure in patients, and the other is the patient’s blood pressure itself. Then, ordinary data analysis can study the existence and size of a stochastic dependency between those two random variables. But if we are interested in the efficiency of the drug in changing blood pressure, this purely stochastic information is insufficient, since there can be a correlation between the two random variables without causation, i.e. without the drug having any effect on blood pressure. Causal analysis on the other hand provides techniques to answer questions about actual causation. And in this example, it answers the question of whether the drug is actually the reason for the change in the patient’s blood pressure and also measures the size of this causal effect. I.e. causal analysis is concerned with the discovery and measurement of actual mechanisms in the underlying system, which in this case would be the patient’s body.
@@
While there are many causal discovery algorithms for the acyclic, causally sufficient situation, very few exist that are also applicable in the more demanding case of cyclic systems with hidden confounders. In this paper, we compare the properties of four of such methods, namely two techniques, ASP-d [11] and ASP-s [4], which are variants of a constraint-based technique using answer set programming, and two variants, LLC-NF and LLC-F [10], of a method of moments type estimator.
@@ II Related Work
For a complete overview of the history of causality see the treatises [15] and [16]. Both references focus mainly on the acyclic case but do cover, to a certain extent, the situation with hidden confounders.
An early technique for cyclic systems, called CCD, is described in [19], but it presumes the absence of latent confounders. Amongst the few algorithms that allow for both cycles and hidden confounders are LLC [10], the method described in [11] which we refer to as ASP-d, sigmasep [4] which we refer to as ASP-s, BACKSHIFT [20], CCI [23], and bcause [17, 18].
Note that the methods above presume interventional data, that is, not only data from the system itself but also from other systems that are obtained by changing the original system in a certain way, i.e. intervening on it. If this interventional data is not available, causal inference becomes harder. There are, however, several approaches for this scenario, too, like the family of Additive Noise Models (ANM), see e.g. [8], ICA-based methods for linear systems like LiNGAM [22], LiNG [12], and the Two-Step algorithm [21], or Information Geometric Causal Inference (IGCI), see e.g. [14]. Some of those methods can also deal with latent confounders or cyclicity.
In recent years, the mathematical foundations of the theory of cyclic systems with hidden confounders have advanced considerably. A comprehensive exposition can be found in [5] and [2].
@@ III Description of the Methods
This section will present the main features of the LLC and ASP algorithms. Note, that this is a high-level overview, presenting only as much as is necessary to explain the evaluation below. For the details, the reader should consult the pertinent papers, see [3, 9, 10] for the LLC and [11, 4] for the ASP variants.
@@ III-A General Concepts in Causality
The main entity in causality is the Structural Causal Model (SCM) [2], which consists of structural equations of the form (note that we use $1\!:\!{n}$ as an abbreviation for the set $\{1,2,\ldots,n\}$):
@@
where $\mathbf{X}=(X_{1},\ldots,X_{n})$ is an $n$-dimensional random vector containing the observed variables, $\mathbf{E}$ is an $m$-dimensional random vector containing the unobserved noise, and the functions $f_{i}:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$ are the causal mechanisms. See [2] for the mathematical details. Many aspects of an SCM can be described by associating a graph that contains as nodes both the observed variables $X_{i},i\in 1\!:\!{n}$ and the unobserved variables $E_{k},k\in 1\!:\!{m}$. It is a directed graph (DG), i.e. all edges have exactly one arrowhead. Here, a directed edge only goes into observed nodes $X_{i}$, no edges are pointing to noise variables $E_{k}$, and there is an edge $N\to X_{i}$ from a node $N$ to $X_{i}$ iff $f_{i}$ depends on $N$. This graph is called the augmented graph of the SCM, see [2].
@@
The directed edges in the graph of an SCM then depict the direct causal connections between variables, symbolizing direct causal effects, while bidirected edges describe confounding (see below). Directed paths consisting of more than one edge describe indirect causal connections, generating indirect causal effects. Note, however, that in cyclic SCMs we can have causal effects even between nodes that are not connected by a directed graph, see [2] section 7.1.
@@
A central notion in causality is that of an intervention. We consider only surgical interventions [10], which can be described as follows: The original SCM is changed by selecting a subset $\mathbf{X}_{I},I\subset 1\!:\!{n}$, of the observed variables, and forcing a new distribution on $\mathbf{X}_{I}$, that is independent of all the other random variables $\mathbf{X}_{1:n\setminus I}$ and $\mathbf{E}$. An example would be randomized drug trials, which ensure that the people who do and do not get the drug are selected completely at random. In this case, we have an intervention on the assignment of the drug. The graph of the intervened SCM differs from the graph of the original SCM in that all the edges that point into intervened nodes are removed.
@@
The idea of constraint-based methods is based on two assumptions (see [15, 16] for details):
@@
    For linear, possibly cyclic SCMs, which is the case we are considering, this assumption holds under mild conditions. See Theorem A.7 in [5] for a precise statement of those conditions.
@@ III-B The LLC algorithm
This section gives a high-level overview of the LLC algorithm; for more details, see [10].
@@
where $\mathbf{x},\mathbf{e}\in\mathbb{R}^{n},\mathbf{B}\in\mathbb{R}^{n\times n}$. Here, $\mathbf{x}$ contains the measurements of the observed variables $\mathbf{X}$ and $\mathbf{e}$ the hidden values of the unobserved variables $\mathbf{E}$. For $\mathbf{E}$, we only presume the expectation to be zero, $\mathbb{E}[\mathbf{E}]=0$, and the covariance of $\mathbf{E}$ is abbreviated with $\mathbf{\Sigma}_{\mathbf{e}}$. Note, that we do not require any particular distribution for $\mathbf{E}$, the only restriction is $\mathbb{E}[\mathbf{E}]=0$, and even that can be lifted, see [10]. However, since this change has no bearing on our evaluation, we stick to the simple case $\mathbb{E}[\mathbf{E}]=0$. The $(i,j)$-coefficient of the matrix $\mathbf{B}$ is denoted by $b_{ij}$ and can be identified as the direct causal effect of variable $X_{j}$ on variable $X_{i}$. $\mathbf{B}$ is allowed to describe cyclic paths in the causal graph. The covariance $\mathbf{\Sigma}_{\mathbf{e}}$ is allowed to have off-diagonal elements, which can be interpreted as confounding. $\mathbf{B}$ is required to have a zero diagonal:
@@
where $\mathbf{t}$ contains all the total effects on the left-hand side of (8), $\mathbf{b}$ is the concatenation of all the rows of $\mathbf{B}$ without the diagonal, i.e. $\mathbf{b}\in\mathbb{R}^{n^{2}-n}$, and $\mathbf{T}$ contains the total effects from the right-hand side of (8). In [10] it is shown, that for $\mathbf{T}$ in (9) to have zero null space, the experiments $\{\mathcal{E}_{k}\}_{k=1}^{K}$ need to satisfy the pair condition, which requires that for each ordered pair of indices $(i,j),i,j\in 1\!:\!{n}$, there is an experiment $\mathcal{E}=(\mathcal{J},\mathcal{U})$ such that $i\in\mathcal{J},j\in\mathcal{U}$.
@@ III-C The LLC-F algorithm
Note, that LLC is not a constraint-based method, and does not presume faithfulness. However, the question is whether combining LLC with constraint-based methods could improve the accuracy of LLC. This has been investigated in [9] and the pertinent algorithm is called LLC-F, with the additional letter “F” indicating that now, faithfulness is presumed. We use the abbreviation LLC-NF to refer to the LLC variant that does not use constraint-based methods and does not presume faithfulness.
@@ III-D ASP-d
Next, we give an overview of the ASP-d algorithm, following [11]. ASP is an abbreviation of “Answer Set Programming”, which is a declarative programming language, see [7].
@@
Several possibilities for how to determine the weights $w(k)$ have been proposed, see e.g. [11, 4, 18]. We will presume that the constraints $k\in\mathbf{K}$ have been obtained using conditional independence hypothesis tests with significance level $\alpha$ and then use weights as in [4]:
@@
The optimization of (13) is done by encoding the problem as an ASP program which can then be solved by some ASP solver like e.g. clingo [6].
@@ III-E ASP-s
In constraint-based methods, the (in)dependences obtained from the measurements must be matched with the graph of the SCM to discover its causal structure. This matching is done, as explained above, via $d$-separation. However, it should be noted that for nonlinear cyclic SCMs, $d$-separation in general fails to be Markovian, see [5].
I.e., for nonlinear cyclic models, $d$-separation must be replaced with another separation property. This new separation property $\sigma$-separation has been introduced in [5]. The basic idea here is roughly that nodes in loops are so strongly connected that conditioning cannot separate them, so they behave as if they would be fully connected. That means, if loops are replaced by fully connected subgraphs, an operation which is called an acyclification, see [5, 2] for details, the resulting acyclic graph exposes $d$-separation properties that again correspond to the conditional (in)dependences of the original cyclic SCM. In [5] the authors then formulated a separation property for cyclic graphs, $\sigma$-separation, which is the “pull back” of $d$-separation via the acyclification operation.
Then, the authors of [4] took the ASP-d algorithm and changed the ASP encoding slightly to now use $\sigma$-separation instead of $d$-separation. In [18] this algorithm got further improved to the ASP-s algorithm we use in this paper.
@@
It is important to note, however, that ASP-d and ASP-s have different application fields. While ASP-d can be used for acyclic and linear cyclic SCMs with hidden confounders, see above, ASP-s applies to acyclic and nonlinear cyclic SCMs with hidden confounders. In particular, linear cyclic SCMs are not faithful w.r.t. $\sigma$-separation, see [5, 4]. Thus, those two application fields have only the acyclic SCMs in common. And since this paper only examines linear SCMs, ASP-s is not strictly applicable here. However, we included ASP-s in the set of algorithms to compare. Since linear models are “of measure zero” inside the set of all SCMs, one might be tempted to always use $\sigma$-separation. This could be problematic, because nonlinear SCMs that are nearly linear could provide data that is only bearly distinguishable from that of linear SCMs. Thus, it would be instructive to see how much worse ASP-s performs compared to ASP-d on linear cyclic models.
@@ IV Evaluation
Here we describe our evaluation of the four methods LLC-NF, LLC-F, ASP-d, and ASP-s. The source code of our implementation in Python and R is available online [^1] The code will be made available upon publication., and uses publicly available code for LLC from Antti Hyttinen’s homepage [^2] https://www.cs.helsinki.fi/u/ajhyttin/ and for ASP from the GitHub repository of [4] [^3] https://github.com/caus-am/sigmasep.
@@
As described in Section III-B, the LLC variants are based on solving the system (9). For real data this system must be expected to contain contradicting equations. This could be handled by solving it as a least squares problem. However, as stated above, the system could also have a non-zero null space if there are not sufficiently diverse experiments to satisfy the pair condition. This would result in the indeterminacy of some or all of the coefficients in $\mathbf{b}$. To avoid this, we chose to use the version of LLC that applies a penalty term ($L_{1}$ or $L_{2}$) to (9), see the code base of [10], which also has the useful effect of regularization and promoting sparsity.
@@
For the evaluation of the methods, we need a metric. Both the LLC and the ASP algorithms return for each feature a score determining how strongly the algorithm “believes” this feature to be present. The code from [10] also provides a bootstrapped version of LLC which means we obtain for each coefficient $b_{ij}$ and $\mathbf{\Sigma}_{\mathbf{e}}$ a collection of estimates of which we can compute the z-score. This is the score we use for both LLC algorithms. For the ASP variants, we utilize a score proposed in [13]: the ASP algorithm is run twice for every single feature, once with the additional constraint that the feature is present and once with the additional constraint that it is absent. The score is then the difference between the loss under the constraint of absence and the loss under the constraint of presence. We call this the ASP confidence score and use this as our scoring function for both ASP variants.
@@
Int. Size
Setup ID
Intervention Sets
0
0
1
11
[], [0]
12
[], [0], [1]
13
[], [0], [1], [2]
14
[], [0], [1], [2], [3]
15
[], [0], [1], [2], [3], [4]
2
21
[], [0,1]
22
[], [0,1], [1,2]
23
[], [0,1], [1,2], [2,3]
24
[], [0,1], [1,2], [2,3], [3,4]
25
[], [0,1], [1,2], [2,3], [3,4], [4,0]
3
31
[], [0,1,2]
32
[], [0,1,2], [1,2,3]
33
[], [0,1,2], [1,2,3], [2,3,4]
34
[], [0,1,2], [1,2,3], [2,3,4], [3,4,0]
35
[], [0,1,2], [1,2,3], [2,3,4], [3,4,0], [4,0,1]
4
41
[], [0,1,2,3]
42
[], [0,1,2,3], [1,2,3,4]
43
[], [0,1,2,3], [1,2,3,4], [2,3,4,0]
44
[], [0,1,2,3], [1,2,3,4], [2,3,4,0], [3,4,0,1]
45
[], [0,1,2,3], [1,2,3,4], [2,3,4,0], [3,4,0,1], [4,0,1,2]
@@
We also consider different sizes $n$ of datasets. We investigate sizes 1000, 10,000, 100,000, and infinite. Here, the word “infinite” does not refer to actually infinite amounts of data, but rather a version of the experiments that correspond to its asymptotic behavior when $n\to\infty$. More precisely, this means the following (see also [10]): When creating random SCMs, the matrices $\mathbf{B}$ and $\mathbf{\Sigma}_{\mathbf{e}}$ are sampled as described above. Then, we can obtain the covariance matrix $\mathbf{C}_{\mathbf{x}}$ of the observations as follows, similar to (10):
@@
The methods we study have certain parameters that need to be fixed. For LLC-NF it is the penalty type, $L_{1}$ or $L_{2}$, and the belonging regularization parameter $\lambda$. For LLC-F we also have to fix the significance level $\alpha_{LLC}$ of the conditional independence tests. Furthermore, when using accuracy as metric, we need to choose a threshold $t_{LLC}$ for the scores. For ASP we have the significance level $\alpha_{ASP}$ for the conditional independence tests and, in the case of using accuracy as a metric, again a threshold $t_{ASP}$ for the score. As with all unsupervised learning tasks, the proper choice of those hyperparameters is tricky in practical applications. Here, we don’t consider the problem of finding optimal hyperparameters but are rather interested in the performance of the studied techniques if the hyperparameters are chosen roughly appropriately, however, this is achieved. Since in our simulations we have access to the ground truth, we can optimize the hyperparameters. For this, we conducted rather comprehensive hyperparameter optimization studies for the type of SCMs that get generated by our adopted sampling scheme which produces sparse, small, cyclic models with hidden confounders as described above. This was done using the optimization tool Optuna [1]. As usual, we used one set of models for hyperparameter optimization and a different one for evaluation. Based on these optimization studies we chose the following hyperparameters for the LLC methods: penalty type $L_{1}$, $\lambda=0.05$, $\alpha_{LLC}=0.05$, and $t_{LLC}=5$. For ASP we used $\alpha_{ASP}=0.05$ and $t_{ASP}=0$.

### v2

# llcsig.bib

@@ Abstract
Nowadays, the need for causal discovery is ubiquitous. A better understanding of not just the stochastic dependencies between parts of a system, but also the actual cause-effect relations, is essential for all parts of science. Thus, the need for reliable methods to detect causal directions is growing constantly. In the last 50 years, many causal discovery algorithms have emerged, but most of them are applicable only under the assumption that the systems have no feedback loops and that they are causally sufficient, i.e. that there are no unmeasured subsystems that can affect multiple measured variables. This is unfortunate since those restrictions can often not be presumed in practice. Feedback is an integral feature of many processes, and real-world systems are rarely completely isolated and fully measured. Fortunately, in recent years, several techniques, that can cope with cyclic, causally insufficient systems, have been developed. And with multiple methods available, a practical application of those algorithms now requires knowledge of the respective strengths and weaknesses. Here, we focus on the problem of causal discovery for sparse linear models which are allowed to have cycles and hidden confounders. We have prepared a comprehensive and thorough comparative study of four causal discovery techniques: two versions of the LLC method [llc] and two variants of the ASP-based algorithm [asp]. The evaluation investigates the performance of those techniques for various experiments with multiple interventional setups and different dataset sizes.
@@ I Introduction
Causal analysis [pearl, pjs] of complex systems is nowadays an integral part of many sciences. It is heavily used in fields as diverse as medicine, biology, cognitive science, economics, predictive maintenance, root cause analysis, physics, and machine learning. Conventional data analysis investigates the probabilistic properties of the data to gain insight into the involved probability distributions which can then be used to e.g. predict new data. Causality on the other hand serves to not just learn the data but to learn about the system that generates the data. For instance, consider two random variables, one is binary and indicates the administration of a drug that is allegedly reducing blood pressure in patients, and the other is the patient’s blood pressure itself. Then, ordinary data analysis can study the existence and size of a stochastic dependency between those two random variables. But if we are interested in the efficiency of the drug in changing blood pressure, this purely stochastic information is insufficient, since there can be a correlation between the two random variables without causation, i.e. without the drug having any effect on blood pressure. Causal analysis on the other hand provides techniques to answer questions about actual causation. And in this example, it answers the question of whether the drug is actually the reason for the change in the patient’s blood pressure and also measures the size of this causal effect. I.e. causal analysis is concerned with the discovery and measurement of actual mechanisms in the underlying system, which in this case would be the patient’s body.
@@
While there are many causal discovery algorithms for the acyclic, causally sufficient situation, very few exist that are also applicable in the more demanding case of cyclic systems with hidden confounders. In this paper, we compare the properties of four of such methods, namely two techniques, ASP-d [asp] and ASP-s [sigmasep], which are variants of a constraint-based technique using answer set programming, and two variants, LLC-NF and LLC-F [llc], of a method of moments type estimator.
@@ II Related Work
For a complete overview of the history of causality see the treatises [pearl] and [pjs]. Both references focus mainly on the acyclic case but do cover, to a certain extent, the situation with hidden confounders.
An early technique for cyclic systems, called CCD, is described in [richardson2013discovery], but it presumes the absence of latent confounders. Amongst the few algorithms that allow for both cycles and hidden confounders are LLC [llc], the method described in [asp] which we refer to as ASP-d, sigmasep [sigmasep] which we refer to as ASP-s, BACKSHIFT [rothenhausler2015backshift], CCI [strobl2019constraint], and bcause [bcause, sBcause].
Note that the methods above presume interventional data, that is, not only data from the system itself but also from other systems that are obtained by changing the original system in a certain way, i.e. intervening on it. If this interventional data is not available, causal inference becomes harder. There are, however, several approaches for this scenario, too, like the family of Additive Noise Models (ANM), see e.g. [hoyer2008nonlinear], ICA-based methods for linear systems like LiNGAM [lingam], LiNG [lacerda2012discovering], and the Two-Step algorithm [twostep], or Information Geometric Causal Inference (IGCI), see e.g. [mooij2016distinguishing]. Some of those methods can also deal with latent confounders or cyclicity.
In recent years, the mathematical foundations of the theory of cyclic systems with hidden confounders have advanced considerably. A comprehensive exposition can be found in [forre2017markov] and [bongers].
@@ III Description of the Methods
This section will present the main features of the LLC and ASP algorithms. Note, that this is a high-level overview, presenting only as much as is necessary to explain the evaluation below. For the details, the reader should consult the pertinent papers, see [eberhardt2010combining, llc-f, llc] for the LLC and [asp, sigmasep] for the ASP variants.
@@ III-A General Concepts in Causality
The main entity in causality is the Structural Causal Model (SCM) [bongers], which consists of structural equations of the form (note that we use $1\!:\!{n}$ as an abbreviation for the set $\{1,2,\ldots,n\}$):
@@
where $\mathbf{X}=(X_{1},\ldots,X_{n})$ is an $n$-dimensional random vector containing the observed variables, $\mathbf{E}$ is an $m$-dimensional random vector containing the unobserved noise, and the functions $f_{i}:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$ are the causal mechanisms. See [bongers] for the mathematical details. Many aspects of an SCM can be described by associating a graph that contains as nodes both the observed variables $X_{i},i\in 1\!:\!{n}$ and the unobserved variables $E_{k},k\in 1\!:\!{m}$. It is a directed graph (DG), i.e. all edges have exactly one arrowhead. Here, a directed edge only goes into observed nodes $X_{i}$, no edges are pointing to noise variables $E_{k}$, and there is an edge $N\to X_{i}$ from a node $N$ to $X_{i}$ iff $f_{i}$ depends on $N$. This graph is called the augmented graph of the SCM, see [bongers].
@@
The directed edges in the graph of an SCM then depict the direct causal connections between variables, symbolizing direct causal effects, while bidirected edges describe confounding (see below). Directed paths consisting of more than one edge describe indirect causal connections, generating indirect causal effects. Note, however, that in cyclic SCMs we can have causal effects even between nodes that are not connected by a directed graph, see [bongers] section 7.1.
@@
A central notion in causality is that of an intervention. We consider only surgical interventions [llc], which can be described as follows: The original SCM is changed by selecting a subset $\mathbf{X}_{I},I\subset 1\!:\!{n}$, of the observed variables, and forcing a new distribution on $\mathbf{X}_{I}$, that is independent of all the other random variables $\mathbf{X}_{1:n\setminus I}$ and $\mathbf{E}$. An example would be randomized drug trials, which ensure that the people who do and do not get the drug are selected completely at random. In this case, we have an intervention on the assignment of the drug. The graph of the intervened SCM differs from the graph of the original SCM in that all the edges that point into intervened nodes are removed.
@@
The idea of constraint-based methods is based on two assumptions (see [pearl, pjs] for details):
@@
    For linear, possibly cyclic SCMs, which is the case we are considering, this assumption holds under mild conditions. See Theorem A.7 in [forre2017markov] for a precise statement of those conditions.
@@ III-B The LLC algorithm
This section gives a high-level overview of the LLC algorithm; for more details, see [llc].
@@
where $\mathbf{x},\mathbf{e}\in\mathbb{R}^{n},\mathbf{B}\in\mathbb{R}^{n\times n}$. Here, $\mathbf{x}$ contains the measurements of the observed variables $\mathbf{X}$ and $\mathbf{e}$ the hidden values of the unobserved variables $\mathbf{E}$. For $\mathbf{E}$, we only presume the expectation to be zero, $\mathbb{E}[\mathbf{E}]=0$, and the covariance of $\mathbf{E}$ is abbreviated with $\mathbf{\Sigma}_{\mathbf{e}}$. Note, that we do not require any particular distribution for $\mathbf{E}$, the only restriction is $\mathbb{E}[\mathbf{E}]=0$, and even that can be lifted, see [llc]. However, since this change has no bearing on our evaluation, we stick to the simple case $\mathbb{E}[\mathbf{E}]=0$. The $(i,j)$-coefficient of the matrix $\mathbf{B}$ is denoted by $b_{ij}$ and can be identified as the direct causal effect of variable $X_{j}$ on variable $X_{i}$. $\mathbf{B}$ is allowed to describe cyclic paths in the causal graph. The covariance $\mathbf{\Sigma}_{\mathbf{e}}$ is allowed to have off-diagonal elements, which can be interpreted as confounding. $\mathbf{B}$ is required to have a zero diagonal:
@@
where $\mathbf{t}$ contains all the total effects on the left-hand side of (8), $\mathbf{b}$ is the concatenation of all the rows of $\mathbf{B}$ without the diagonal, i.e. $\mathbf{b}\in\mathbb{R}^{n^{2}-n}$, and $\mathbf{T}$ contains the total effects from the right-hand side of (8). In [llc] it is shown, that for $\mathbf{T}$ in (9) to have zero null space, the experiments $\{\mathcal{E}_{k}\}_{k=1}^{K}$ need to satisfy the pair condition, which requires that for each ordered pair of indices $(i,j),i,j\in 1\!:\!{n}$, there is an experiment $\mathcal{E}=(\mathcal{J},\mathcal{U})$ such that $i\in\mathcal{J},j\in\mathcal{U}$.
@@ III-C The LLC-F algorithm
Note, that LLC is not a constraint-based method, and does not presume faithfulness. However, the question is whether combining LLC with constraint-based methods could improve the accuracy of LLC. This has been investigated in [llc-f] and the pertinent algorithm is called LLC-F, with the additional letter “F” indicating that now, faithfulness is presumed. We use the abbreviation LLC-NF to refer to the LLC variant that does not use constraint-based methods and does not presume faithfulness.
@@ III-D ASP-d
Next, we give an overview of the ASP-d algorithm, following [asp]. ASP is an abbreviation of “Answer Set Programming”, which is a declarative programming language, see [aspLogic].
@@
Several possibilities for how to determine the weights $w(k)$ have been proposed, see e.g. [asp, sigmasep, sBcause]. We will presume that the constraints $k\in\mathbf{K}$ have been obtained using conditional independence hypothesis tests with significance level $\alpha$ and then use weights as in [sigmasep]:
@@
The optimization of (13) is done by encoding the problem as an ASP program which can then be solved by some ASP solver like e.g. clingo [clingo].
@@ III-E ASP-s
In constraint-based methods, the (in)dependences obtained from the measurements must be matched with the graph of the SCM to discover its causal structure. This matching is done, as explained above, via $d$-separation. However, it should be noted that for nonlinear cyclic SCMs, $d$-separation in general fails to be Markovian, see [forre2017markov].
I.e., for nonlinear cyclic models, $d$-separation must be replaced with another separation property. This new separation property $\sigma$-separation has been introduced in [forre2017markov]. The basic idea here is roughly that nodes in loops are so strongly connected that conditioning cannot separate them, so they behave as if they would be fully connected. That means, if loops are replaced by fully connected subgraphs, an operation which is called an acyclification, see [forre2017markov, bongers] for details, the resulting acyclic graph exposes $d$-separation properties that again correspond to the conditional (in)dependences of the original cyclic SCM. In [forre2017markov] the authors then formulated a separation property for cyclic graphs, $\sigma$-separation, which is the “pull back” of $d$-separation via the acyclification operation.
Then, the authors of [sigmasep] took the ASP-d algorithm and changed the ASP encoding slightly to now use $\sigma$-separation instead of $d$-separation. In [sBcause] this algorithm got further improved to the ASP-s algorithm we use in this paper.
@@
It is important to note, however, that ASP-d and ASP-s have different application fields. While ASP-d can be used for acyclic and linear cyclic SCMs with hidden confounders, see above, ASP-s applies to acyclic and nonlinear cyclic SCMs with hidden confounders. In particular, linear cyclic SCMs are not faithful w.r.t. $\sigma$-separation, see [forre2017markov, sigmasep]. Thus, those two application fields have only the acyclic SCMs in common. And since this paper only examines linear SCMs, ASP-s is not strictly applicable here. However, we included ASP-s in the set of algorithms to compare. Since linear models are “of measure zero” inside the set of all SCMs, one might be tempted to always use $\sigma$-separation. This could be problematic, because nonlinear SCMs that are nearly linear could provide data that is only bearly distinguishable from that of linear SCMs. Thus, it would be instructive to see how much worse ASP-s performs compared to ASP-d on linear cyclic models.
@@ IV Evaluation
Here we describe our evaluation of the four methods LLC-NF, LLC-F, ASP-d, and ASP-s. The source code of our implementation in Python and R is available online [^1] https://github.com/mustafa-mohsen/llcsig, and uses publicly available code for LLC from Antti Hyttinen’s homepage [^2] https://www.cs.helsinki.fi/u/ajhyttin/ and for ASP from the GitHub repository of [sigmasep] [^3] https://github.com/caus-am/sigmasep.
@@
As described in Section III-B, the LLC variants are based on solving the system (9). For real data this system must be expected to contain contradicting equations. This could be handled by solving it as a least squares problem. However, as stated above, the system could also have a non-zero null space if there are not sufficiently diverse experiments to satisfy the pair condition. This would result in the indeterminacy of some or all of the coefficients in $\mathbf{b}$. To avoid this, we chose to use the version of LLC that applies a penalty term ($L_{1}$ or $L_{2}$) to (9), see the code base of [llc], which also has the useful effect of regularization and promoting sparsity.
@@
For the evaluation of the methods, we need a metric. Both the LLC and the ASP algorithms return for each feature a score determining how strongly the algorithm “believes” this feature to be present. The code from [llc] also provides a bootstrapped version of LLC which means we obtain for each coefficient $b_{ij}$ and $\mathbf{\Sigma}_{\mathbf{e}}$ a collection of estimates of which we can compute the z-score. This is the score we use for both LLC algorithms. For the ASP variants, we utilize a score proposed in [magliacane2016ancestral]: the ASP algorithm is run twice for every single feature, once with the additional constraint that the feature is present and once with the additional constraint that it is absent. The score is then the difference between the loss under the constraint of absence and the loss under the constraint of presence. We call this the ASP confidence score and use this as our scoring function for both ASP variants.
@@
Int. Size Setup ID Intervention Sets
0 0 []
1 11 [], [0]
12 [], [0], [1]
13 [], [0], [1], [2]
14 [], [0], [1], [2], [3]
15 [], [0], [1], [2], [3], [4]
2 21 [], [0,1]
22 [], [0,1], [1,2]
23 [], [0,1], [1,2], [2,3]
24 [], [0,1], [1,2], [2,3], [3,4]
25 [], [0,1], [1,2], [2,3], [3,4], [4,0]
3 31 [], [0,1,2]
32 [], [0,1,2], [1,2,3]
33 [], [0,1,2], [1,2,3], [2,3,4]
34 [], [0,1,2], [1,2,3], [2,3,4], [3,4,0]
35 [], [0,1,2], [1,2,3], [2,3,4], [3,4,0], [4,0,1]
4 41 [], [0,1,2,3]
42 [], [0,1,2,3], [1,2,3,4]
43 [], [0,1,2,3], [1,2,3,4], [2,3,4,0]
44 [], [0,1,2,3], [1,2,3,4], [2,3,4,0], [3,4,0,1]
45 [], [0,1,2,3], [1,2,3,4], [2,3,4,0], [3,4,0,1], [4,0,1,2]
@@
We also consider different sizes $n$ of datasets. We investigate sizes 1000, 10,000, 100,000, and infinite. Here, the word “infinite” does not refer to actually infinite amounts of data, but rather a version of the experiments that correspond to its asymptotic behavior when $n\to\infty$. More precisely, this means the following (see also [llc]): When creating random SCMs, the matrices $\mathbf{B}$ and $\mathbf{\Sigma}_{\mathbf{e}}$ are sampled as described above. Then, we can obtain the covariance matrix $\mathbf{C}_{\mathbf{x}}$ of the observations as follows, similar to (10):
@@
The methods we study have certain parameters that need to be fixed. For LLC-NF it is the penalty type, $L_{1}$ or $L_{2}$, and the belonging regularization parameter $\lambda$. For LLC-F we also have to fix the significance level $\alpha_{LLC}$ of the conditional independence tests. Furthermore, when using accuracy as metric, we need to choose a threshold $t_{LLC}$ for the scores. For ASP we have the significance level $\alpha_{ASP}$ for the conditional independence tests and, in the case of using accuracy as a metric, again a threshold $t_{ASP}$ for the score. As with all unsupervised learning tasks, the proper choice of those hyperparameters is tricky in practical applications. Here, we don’t consider the problem of finding optimal hyperparameters but are rather interested in the performance of the studied techniques if the hyperparameters are chosen roughly appropriately, however, this is achieved. Since in our simulations we have access to the ground truth, we can optimize the hyperparameters. For this, we conducted rather comprehensive hyperparameter optimization studies for the type of SCMs that get generated by our adopted sampling scheme which produces sparse, small, cyclic models with hidden confounders as described above. This was done using the optimization tool Optuna [optuna_2019]. As usual, we used one set of models for hyperparameter optimization and a different one for evaluation. Based on these optimization studies we chose the following hyperparameters for the LLC methods: penalty type $L_{1}$, $\lambda=0.05$, $\alpha_{LLC}=0.05$, and $t_{LLC}=5$. For ASP we used $\alpha_{ASP}=0.05$ and $t_{ASP}=0$.